In a particular species of plant, tall is dominant to short, and orange petals are dominant to the recessive white colour. Use T and t to symbolize the alleles for height, and F and f to symbolize the alleles for flower colour. A homozygous tall white flower is crossed with a flower heterozygous for both traits.

1.1 List the genotypes of the parents.                                           

1.2 What are the F1 GENOTYPIC and PHENOTYPIC ratios?          [5 marks]

1.3 What are the F2 genotypic and phenotypic ratios?   

 


1
Expert's answer
2021-09-27T17:01:01-0400

a)

The parents phenotype are

TTff x TtFf

b)

The f1 genotype will be

TTFf, TTff, TtFf, Ttff

The phenotypic ratio of tall and orange colored is 4/8= ½

The phenotypic ratio for tall and white is 4/8= ½

The genotypic ratio for each is ¼

c)

The f1 genotype will be

The phenotypic ratio of tall and orange colored is 4/8= ½

The phenotypic ratio for tall and white is 4/8= ½

The genotypic ratio for each is 3/4